# A Comprehensive Reanalysis of K2-18 b’s JWST NIRISS+NIRSpec Transmission Spectrum

### v1

@@ 4.1.2 BEAR
To validate our POSEIDON results, we also perform a second set of retrievals using the open-source Bern Atmospheric Retrieval code (BeAR888https://newstrangeworlds.github.io/BeAR). BeAR is an updated version of the retrieval code formerly known as Helios-r2, which was originally developed for retrievals on Brown Dwarf spectra (Kitzmann et al., 2020). BeAR has been adapted for use on exoplanet spectra (Fisher et al., 2024) and has the capabilities to analyze transmission, emission, and phase curve spectra. BeAR uses opacity sampling of line-by-line cross sections to calculate model spectra, coupled to PyMultiNest (Feroz et al., 2009; Buchner et al., 2014) for statistical exploration.
@@ 4.2.2 Molecular Abundance Constraints
Next, we consider the molecular abundance constraints from our full collection of POSEIDON retrievals. For this purpose, we seek a way to combine the posterior distributions corresponding to each of the 60 data combinations. Given the identical parameter space of each of the 60 reference model retrievals, we can combine the samples from each retrieval to produce a single ensemble posterior. This technique essentially marginalizes over two additional dimensions, NIRISS and NIRSpec data-level choices, and thereby folds the impacts of three different data reduction codes and different data resolutions into a single set of atmospheric abundance constraints. To our knowledge, this conceptually simple — but computationally expensive — technique is novel for atmospheric inferences from JWST data. We present our ensemble posterior for K2-18 b’s atmospheric composition, spanning over 30 million likelihood evaluations, in Figure 7.
@@
CH4 is the only molecule with an abundance lower limit — indicating a robust detection — in Figure 7. Our ensemble posterior finds a well-constrained abundance of , which is consistent with the value of reported in Madhusudhan et al. (2023) (see also Appendix A). This high abundance of CH4 in K2-18 b’s atmosphere, % by volume, is consistent with a metal-enriched mini-Neptune (as discussed in Section 5.1) and is similar to the deep abundance of CH4 on Neptune (%; Karkoschka & Tomasko 2011).
We report upper limits on the abundances of CO2 and DMS in K2-18 b’s atmosphere: CO and DMS . The ensemble posteriors for these molecules in Figure 7 show a weak peak for DMS, alongside an even more marginal peak for CO2, which reflects the combination of retrieval results from different data combinations where a few favor including DMS or CO2 but most do not (e.g. Figure 6, top panel). The blending of possible CO2 and DMS posteriors from all our data variants into a single conservative posterior, guarding against statistical artifacts from specific data-level assumptions (e.g. Figure 5), is a distinct strength of our ensemble posterior method. We stress that these upper limits do not imply that either molecule is absent at any atmospheric abundance, as abundance lower than our 95% limits are still compatible with the present JWST transmission spectrum. Indeed, some mini-Neptune scenarios predict CO2 abundances of (Section 5.1), consistent with our upper limit. Consequently, our non-detection of CO2 with these data does not preclude more sensitive upcoming data (e.g. JWST Program GO-2372, PI: Renyu Hu) from eventually identifying CO2 and other trace species in K2-18 b’s atmosphere.
@@
We additionally report upper limits on several gases important for understanding the atmospheres and interiors of sub-Neptunes. Water and ammonia are depleted from the gas phase (95% upper limits of and ), which can be explained by cloud condensation (Hu et al., 2021; Madhusudhan et al., 2023) and dissolution into a deep magma ocean (Shorttle et al., 2024), respectively. We show in Section 5 that neither of these non-detections requires a surface liquid ocean on K2-18 b. Alongside our non-detection of DMS, we find upper limits on four other gases proposed as biosignatures (e.g., Segura et al., 2005; Seager et al., 2013; Schwieterman et al., 2018; Madhusudhan et al., 2021): CS, CH3Cl , OCS , and N2O . Therefore, there is no evidence of biosignatures in the JWST NIRISS SOSS + NIRSpec G395H transmission spectrum of K2-18 b.
@@ 4.2.3 Atmospheric Temperature
K2-18 b’s transmission spectrum is consistent with an isothermal upper-atmosphere temperature structure. While our POSEIDON retrievals allow for temperature gradients, our top-of-atmosphere temperature ( K) is consistent with the temperature at significantly higher pressures (e.g. = K, see Figure 7). Our temperature constraints are consistent with Madhusudhan et al. (2023), who reported K. Since low-resolution transmission spectra only probe atmospheric pressures from – bar, we note that the lack of any detected temperature gradient does not preclude steep adiabatic temperature gradients at deeper pressures.
@@ 4.2.4 Aerosols
We do not detect any evidence of aerosols in K2-18 b’s atmosphere. Figure 7 highlights that any optically-thick cloud deck must reside deep in the atmosphere, but the present JWST data can only place a very loose lower limit (). Our retrievals offer no constraints (either upper or lower limits) on the haze parameters or the terminator cloud fraction. Madhusudhan et al. (2023) similarly reported a non-detection of aerosols. Since scattering hazes can play an important role in moderating K2-18 b’s temperature under a hycean interpretation (Madhusudhan et al., 2023), our non-detection of a haze — based on data that include the bluer 0.6–0.8 2nd order NIRISS SOSS data that are in theory more sensitive to a haze slope — provides important context for the viability of a liquid water ocean. However, given the present signal-to-noise ratio at the shortest wavelengths (e.g. Figure 6), additional observations with NIRISS SOSS would be required before hazes can be definitively ruled out.
4.2.5 Planetary Mass
Our retrieval analysis also offers an updated constraint on K2-18 b’s planetary mass. Transmission spectra can, in principle, constrain an exoplanet’s mass through the amplitude and shape of absorption features (de Wit & Seager, 2013), with the tightest constraints arising for cloud-free atmospheres (Batalha et al., 2019a). Given the lack of detected aerosols in K2-18 b’s transmission spectrum, this technique provides the opportunity to potentially improve the mass determination from radial velocity measurements (Cloutier et al., 2019; Radica et al., 2022a). Our POSEIDON retrievals adopt a Gaussian prior on the planetary mass to reflect our existing knowledge from radial velocities (; Cloutier et al. 2019), such that any additional information in the JWST spectrum would result in a tighter constraint. Our ensemble posterior (Figure 7) offers a new constraint on K2-18 b’s mass, , which slightly decreases the mass uncertainty from 16% to 13%. This updated mass determination can be used to inform interior structure models of K2-18 b.
@@ 4.2.6 Unocculted Starspots
By including stellar contamination in our retrievals, we obtain constraints of unocculted active regions on K2-18. Our ensemble posterior (Figure 7) generally indicates a small coverage fraction (%) of cool spots ( K cooler than the photosphere). While this may appear contrary to the finding of no stellar contamination reported by Madhusudhan et al. (2023), we note that our inclusion of the 2nd order NIRISS data provides coverage of shorter wavelengths where starspot contamination is more prominent (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024a; Cadieux et al., 2024; Radica et al., 2024a). We investigated the impact of stellar contamination by running several additional POSEIDON retrievals with stellar contamination excluded, with one example shown in Figure 8. We find that this potential low-level stellar contamination does not notably affect the inferred atmospheric composition of K2-18 b. However, given the stochastic nature of the transit light source effect, further repeat observations will be necessary to confirm this. Therefore, while it is likely that future transit observations of K2-18 b will be able to be stacked to increase the transmission spectrum precision, it would be prudent to continue observations with NIRISS SOSS (as opposed to only observing at longer wavelengths) to further constrain the possibility of stellar heterogeneities.
@@ 5.2 Interior Structure Modeling
To aid in evaluating the possibility of a liquid water ocean on K2-18 b, we also construct several interior structure models of the planet which match its observed mass and radius (, ). We use the mass and radius presented by Benneke et al. (2019) as the basis of our models101010If the planet were smaller, as reported in earlier work, the required water fraction for a thin H/He atmosphere drops to a more plausible range.. We prefer this set of planet parameters due to the use of a Gaia parallax in their calculation, though we note that this mass is also consistent with the mass recovered from our atmosphere retrievals (Section 4.2.5). These interior models typically solve the equations of hydrostatic equilibrium, mass conservation, and the material equations of state. We used the Chabrier & Debras (2021) equation of state for H/He, the Mazevet et al. (2019) equation of state for water, ANEOS (Thompson, 1990) for rock, and SESAME (Lyon & Johnson, 1992) for iron, and thermally evolve the planet using atmosphere models from Fortney et al. (2007) to regulate heat flow from the interior.
@@
We ran an MCMC to find compositions of the planet which match the planet’s observed properties, with the radius being of particular importance. Figure 13 shows the results as a ternary diagram of the mass fractions of rock (and iron, assuming an Earth-like ratio), water, and H/He (at a solar ratio). Our models find H/He makes up less than 10% of the planet’s mass, with the exact limit varying with the water abundance. Any of the models along the shaded region in Figure 13 fits the observed data, but we identify a few specific cases of interest (denoted with circles) for further discussion.
@@
First, we consider the ice-poor case, which represents a typical Neptune-like planet that could form within the ice line, experiencing limited volatile accretion. In this case, approximately 6% of the mass must be H/He to match the observed radius. Such a planet features a deep envelope (due to the low gravity) on top of a solid (possibly molten) surface. If water is mixed into the envelope, it could result in water clouds and rain in the atmosphere, but the envelope is too deep to allow for a water ocean—any falling rain would pass into the supercritical phase and dissolve back into the ambient gas.
We next consider a more ice-rich structure, similar to outer solar system moons like Ganymede, with a 50-50 rock-to-water mass ratio and a structure that is stratified into separate layers (iron, rock, water, H/He). Such a planet requires a H/He envelope of about 1% of the planet’s mass. This results in the surface of the water layer occurring at about 2.21 R⊕, a pressure of 100 kbar and a temperature of 3200 K. This temperature and pressure are high enough that the water is a supercritical fluid rather than a liquid phase.
To find a model which allows for a liquid water ocean, we must consider cases where the H/He layer is very thin. This possibility occurs only at very high water fractions, at least 70%, beyond what one would expect from planetary formation models. In this case, the water layer may extend to temperatures below the critical point of 647 K. However, the H/He adiabats produced by our evolution model reach this temperature at 220 bar, approximately passing through the critical point; thus, the H/He layer must not pressurize the surface to more than 220 bars. This corresponds with a H/He layer mass fraction of no more than about 0.03%. The stability of such an arrangement would depend strongly on the tendency of H/He to dissolve in water or the water to evaporate into the atmosphere (e.g. Pierrehumbert, 2023), but that is outside the scope of this work.
Overall, we conclude that the most likely scenarios are the water-poor and Ganymede-like 50-50 composition models, neither of which features a water ocean. While we are able to construct models that feature a water ocean via heavy water enrichment, it is not clear how planetary formation could achieve such a high water fraction, nor whether the correspondingly thin H/He atmosphere would be stable. Given our non-detection of CO2 already disfavors the hycean model for K2-18 b (as discussed above in Section 5.1), our findings from these interior structure models are further evidence that K2-18 b’s structure is more Neptune-like in nature.
@@ 6 Summary & Discussion
Any inferences of CO2 and DMS from these JWST data are unreliable and low-significance, exhibiting high sensitivity to specific data reduction codes and spectral resolutions (Figure 5).
We present a simple technique to account for data-level choices across a grid of atmospheric retrievals. The resulting ‘ensemble posterior’ folds uncertainties from data processing into atmospheric inferences (Figure 7).
@@
6.1 Plausibility of Hycean Scenarios for K2-18 b
The debate surrounding K2-18 b’s nature since the initial JWST transmission spectrum was published has relied on the abundance constraints reported by Madhusudhan et al. (2023). Shorttle et al. (2024), Yang & Hu (2024), and Wogan et al. (2024) all showed that variations on mini-Neptune scenarios can explain the abundances from Madhusudhan et al. (2023). Wogan et al. (2024) additionally showed that an ‘inhabited’ hycean scenario (with a large CH4 surface flux) is consistent with the chemical abundances from Madhusudhan et al. (2023), but favored the mini-Neptune scenario based on the fewer required assumptions. Recently, Cooke & Madhusudhan (2024) argued that a mini-Neptune scenario is “not compatible” with the retrieved abundances from Madhusudhan et al. (2023), but either lifeless or inhabited hycean scenarios are viable (depending on specific photochemical model assumptions). We argue here that much of this debate stems from taking the abundance constraints and upper limits from Madhusudhan et al. (2023) as the ‘ground truth’ composition of K2-18 b. Since the original retrieval analysis of Madhusudhan et al. (2023) did not explore the sensitivity of the atmospheric composition to data-level or retrieval-level assumptions, the abundances quoted in that study are not sufficiently reliable for discrimination of whether K2-18 b is a mini-Neptune or a hycean planet.
@@ 6.1 Plausibility of Hycean Scenarios for K2-18 b
Our revised atmospheric composition of K2-18 b, accounting for data-level uncertainties, can be naturally explained by a mini-Neptune without requiring a liquid water ocean or life. Our non-detection of CO2 lowers the potential CO2/CH4 from to 0.3 (95% confidence), whereas a high CO2/CH4 is one of the primary lines of evidence for a liquid water ocean underlying a thin H2-dominated envelope (Hu et al., 2021; Tsai et al., 2021a). However, absence of evidence is not evidence of absence, and we stress that even mini-Neptune scenarios will contain some level of CO2 (e.g., Figure 9 predicts ). Consequently, should future, more sensitive, JWST observations result in a CO2 detection with an abundance lower than our upper limit (), this would be an additional line of evidence against a hycean interpretation (see also Section 6.4). Our photochemical modeling offers an additional line of evidence against inhabited hycean scenarios. We show in Figure 12 that, even if one assumes a hycean scenario with a large unknown CH4 flux (e.g. from methane-producing life), over time the atmospheric CO2 and CO abundances would build to levels comparable to or above our upper limits. Consequently, our non-detections of CO2 and CO are not consistent with the observed CH4 abundance under an inhabited hycean scenario. Our results, therefore, render the mini-Neptune scenario the most likely interpretation for K2-18 b, given current observational constraints.
6.2 Viability of a Biosphere on K2-18 b
The potential evidence for a biosphere on K2-18 b has revolved around the large CH4 abundance, requiring a significant surface flux under a hycean scenario, and the claimed spectral signature of the biosignature candidate DMS (Madhusudhan et al., 2023). Glein (2024) showed that metabolic reactions consuming CO2 and/or DMS and producing CH4 can potentially produce sufficient free energy to sustain a biosphere under a hycean scenario on K2-18 b. Glein (2024) highlighted that DMS should not strictly be considered a biosignature under this metabolic pathway, as DMS would be an energy source for methane-producing life rather than a product of life. The calculations from Glein (2024) took the retrieved abundances from Madhusudhan et al. (2023) as the present-day atmospheric composition and assumed K2-18 b is a hycean world.
@@ 6.2 Viability of a Biosphere on K2-18 b
Our revised atmospheric composition motivates a careful reassessment of the viability of methanogenesis on K2-18 b. Our non-detections of CO2 and DMS prompt a re-assessment of whether a sufficient free energy source exists for methane-producing life under the metabolic pathway studied by Glein (2024), rendering it less plausible that K2-18 b sustains a biosphere. Further, our models in Figure 12 demonstrate that a hycean scenario with the CH4 abundance we retrieve would necessarily have a high 1% CO2 abundance that is in tension with our non-detection of CO2. Updated metabolic calculations, considering the minimum viable CO2 abundance to sustain methanogenesis, will be required to assess whether a biosphere is ruled out by our revised atmospheric composition.

### v2

@@ 4.1.2 BEAR
To validate our POSEIDON results, we also perform a second set of retrievals using the open-source Bern Atmospheric Retrieval code (BeAR999https://newstrangeworlds.github.io/BeAR). BeAR is an updated version of the retrieval code formerly known as Helios-r2, which was originally developed for retrievals on Brown Dwarf spectra (Kitzmann et al., 2020). BeAR has been adapted for use on exoplanet spectra (Fisher et al., 2024) and has the capabilities to analyze transmission, emission, and phase curve spectra. BeAR uses opacity sampling of line-by-line cross sections to calculate model spectra, coupled to PyMultiNest (Feroz et al., 2009; Buchner et al., 2014) for statistical exploration.
@@ 4.2.2 Molecular Abundance Constraints
CH4 is the only molecule with an abundance lower limit across all data combinations (Figure 7), which indicates a robust detection. To account for the spread in abundances across different data treatments, we consider a conservative range spanning the lowest and highest across all the CH4 posteriors — similar to the approach by Powell et al. (2024) for SO2 in WASP-39b. We find a range of -2.14 to -0.53 and a range of -2.86 to -0.35, with the most probable CH4 abundance near 10%. Our retrieved CH4 is consistent with the value reported in Madhusudhan et al. (2023) using a single data reduction code (; see also Appendix A). The high abundance of CH4 in K2-18 b’s atmosphere is consistent with a metal-enriched mini-Neptune (as discussed in Section 5.1) and is similar to the deep abundance of CH4 on Neptune (%; Karkoschka & Tomasko 2011). This high CH4 abundance also raises the background mean molecular weight of K2-18 b’s atmosphere to 2.46–7.64 amu.
Abundance inferences for CO2 and DMS depend critically on the data combination adopted, with some combinations showing suggestive peaks but most showing broad posteriors with no evidence supporting their presence (Figure 7). We therefore report upper limits on the abundances of CO2 and DMS in K2-18 b’s atmosphere, since neither of these molecules are robustly detected. Considering extrema upper limits, as for CH4 above, we find CO and DMS . However, while these limits are pulled to higher abundances by the few data combinations exhibiting posterior peaks (Figure 7). Therefore, we consider the median upper limits to provide a more representative statistic for the abundances of molecules that are not robustly detected: CO and DMS . We stress that these upper limits do not imply that either molecule is absent at any atmospheric abundance, as abundances lower than our 95% limits are still compatible with the present JWST transmission spectrum. Indeed, some mini-Neptune scenarios predict CO2 abundances of (Section 5.1), consistent with our upper limit.
@@
We additionally report upper limits on several gases important for understanding the atmospheres and interiors of sub-Neptunes. Water and ammonia are depleted from the gas phase (95% upper limits of and ), which can be explained by cloud condensation (Hu et al., 2021; Madhusudhan et al., 2023) and dissolution into a deep magma ocean (Shorttle et al., 2024), respectively. Similarly, we find an upper limit of . We show in Section 5 that these non-detections do not require a surface liquid ocean on K2-18 b. Alongside our non-detection of DMS, we find upper limits on four other gases proposed as biosignatures (e.g., Segura et al., 2005; Seager et al., 2013; Schwieterman et al., 2018; Madhusudhan et al., 2021): CS, CH3Cl , OCS , and N2O . Therefore, there is no evidence of biosignatures in the JWST NIRISS SOSS + NIRSpec G395H transmission spectrum of K2-18 b.
@@ 4.2.3 Atmospheric Temperature
K2-18 b’s transmission spectrum is consistent with an isothermal upper-atmosphere temperature structure. While our POSEIDON retrievals allow for temperature gradients, our top-of-atmosphere temperature ( 98–339 K) is consistent with the temperature at significantly higher pressures (e.g. = 120–358 K). Our temperature constraints are consistent with Madhusudhan et al. (2023), who reported K. Since low-resolution transmission spectra only probe atmospheric pressures from – bar, we note that the lack of any detected temperature gradient does not preclude steep adiabatic temperature gradients at deeper pressures.
@@ 4.2.4 Aerosols
We do not detect any evidence of aerosols in K2-18 b’s atmosphere. Our retrievals offer no constraints (either upper or lower limits) on the haze parameters or the terminator cloud fraction. Madhusudhan et al. (2023) similarly reported a non-detection of aerosols. Since scattering hazes can play an important role in moderating K2-18 b’s temperature under a hycean interpretation (Madhusudhan et al., 2023; Jordan et al., 2025), our non-detection of a haze — based on data that include the bluer 0.6–0.8 2nd order NIRISS SOSS data that are more sensitive to a haze slope — provides important context for the viability of a liquid water ocean. However, given the present signal-to-noise ratio at the shortest wavelengths (e.g. Figure 6), additional observations with NIRISS SOSS would be required before hazes can be definitively ruled out.
@@ 4.2.5 Unocculted Starspots
By including stellar contamination in our retrievals, we obtain constraints of unocculted active regions on K2-18. We find a small coverage fraction (2–10%) of cool spots ( K cooler than the photosphere). While this may appear contrary to the finding of no stellar contamination reported by Madhusudhan et al. (2023), we note that our inclusion of the 2nd order NIRISS data provides coverage of shorter wavelengths where starspot contamination is more prominent (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024a; Cadieux et al., 2024; Radica et al., 2024a). We investigated the impact of stellar contamination by running several additional POSEIDON retrievals with stellar contamination excluded, with one example shown in Figure 8. We find that this potential low-level stellar contamination does not notably affect the inferred atmospheric composition of K2-18 b. However, given the stochastic nature of the transit light source effect, further repeat observations will be necessary to confirm this. Therefore, while it is likely that future transit observations of K2-18 b will be able to be stacked to increase the transmission spectrum precision, it would be prudent to continue observations with NIRISS SOSS (as opposed to only observing at longer wavelengths) to further constrain the possibility of stellar heterogeneities.
@@ 5.2 Interior Structure Modeling
To aid in evaluating the possibility of a liquid water ocean on K2-18 b, we also construct several interior structure models of the planet which match its observed mass and radius (, ). We use the mass and radius presented by Benneke et al. (2019a) as the basis of our models111111If the planet were smaller, as reported in earlier work, the required water fraction for a thin H/He atmosphere drops to a more plausible range.. We prefer this set of planet parameters due to the use of a Gaia parallax in their calculation. These interior models solve the equations of hydrostatic equilibrium, mass conservation, and the material equations of state (see Thorngren et al., 2016). This is similar to the analysis of Madhusudhan et al. (2020), but includes methane in the atmosphere to comport with the observed spectra, setting its mass fraction prior as normally distributed with mean 0.3 and standard deviation 0.05. We used the Chabrier & Debras (2021) equation of state for H/He, Mazevet et al. (2019) for water, Nettelmann et al. (2016) and Bethkenhagen et al. (2017) for methane, ANEOS (Thompson, 1990) for rock, and SESAME (Lyon & Johnson, 1992) for iron. The methane and H/He are mixed in the atmosphere via the additive volumes approximation, with the adiabats calculated through thermal integration (see Thorngren, 2019, Eqs. 1.07-1.18). We thermally evolve the planet using atmosphere models from Fortney et al. (2007) to regulate heat flow from the interior.
@@
We used MCMC to find compositions of the planet which match the planet’s observed properties, especially radius and the atmospheric methane abundance. Figure 13 shows the results as a ternary diagram of the mass fractions of rock (and iron, assuming an Earth-like ratio), nebular ices (here, water and methane), and H/He (at a solar ratio) for uniformly-distributed composition priors. Our models find H/He makes up less than 10% of the planet’s mass, with the exact limit varying with the water abundance. Any of the models along the shaded region in Figure 13 fits the observed data, but we identify a few specific cases of interest (denoted with circles) for further discussion.
@@
First, we consider the ice-poor case, which represents a typical Neptune-like planet that could form within the ice line, experiencing limited volatile accretion. In this case, approximately 6% of the mass must be H/He to match the observed radius. Such a planet features a deep envelope on top of a solid or molten rock surface. A more ice-rich planet could result from forming outside the ice line, with rock-to-water ratios around 50-50 as seen in the solar system icy moons (Showman & Malhotra, 1999; Kuskov & Kronrod, 2005). Stratified into separate layers, the planet would have a H/He envelope of about 1% of the planet’s mass. This results in the surface of the water layer occurring at about 2.21 R⊕, a pressure of 100 kbar and a temperature of 3200 K; a supercritical fluid rather than a liquid.
To support a liquid water ocean, the surface must not exceed the critical point of water. Our model atmospheres pass almost directly through the water critical point at 647 K and 220 bar. Thus the H/He atmosphere must not pressurize the water beyond this, which corresponds to an H/He layer mass fraction of no more than 0.03%. Our posterior probability of this composition is just 0.59% (Fig. 13, inset). Furthermore, accounting for the possibility of a runaway greenhouse effect could push this pressure limit down to tens of bars (Innes et al., 2023).
In order to match the observed radius, such a low H/He fraction must be compensated for by a high water fraction of at least 70%. This is improbable from a formation perspective – (Lodders, 2003) find the solar water-to-rock ratio to be 1.17:1 (54%), similar to the solar system icy moons (Kuskov & Kronrod, 2005). Marcus et al. (2010) argue that this initial water-to-rock ratio provides an upper limit for the final planet composition, as giant impacts cannot enhance the water fraction, only deplete it. Including other ices (methane and ammonia) gives a max ice-to-rock ratio of 2.41:1 (70%), but this would require formation outside the methane ice line and subsequent migration, and methane’s low solubility in water suggests it would likely contribute to a thick atmosphere rather than permit a water ocean. Thus, while we cannot absolutely rule out a water ocean, such a structure is very improbable for K2-18 b due to its large radius.
@@ 6 Summary & Discussion
Any inferences of CO2 and DMS from these JWST data are unreliable and low-significance, exhibiting high sensitivity to choices made during the reduction process and retrieval model assumptions (Figure 5).
@@
6.1 Plausibility of Hycean Scenarios & Biosphere Viability for K2-18 b
@@
Our revised atmospheric composition of K2-18 b, accounting for data-level uncertainties, can be naturally explained by a mini-Neptune without requiring a liquid water ocean or life. Our non-detection of CO2 lowers the potential CO2/CH4 ratio from to 0.3 (at 95% confidence). This lower ratio complicates a hycean interpretation for K2-18 b, as a high CO2/CH4 is one of the primary lines of evidence for a liquid water ocean underlying a thin H2-dominated envelope (Hu et al., 2021; Tsai et al., 2021a). However, absence of evidence is not evidence of absence, and we stress that even mini-Neptune scenarios will contain some level of CO2 (e.g., Figure 9 predicts ). Consequently, should future, more sensitive, JWST observations result in a CO2 detection with an abundance lower than our upper limit (), this would be an additional line of evidence against a hycean interpretation (see also Section 6.4). Our photochemical modeling offers an additional line of evidence against inhabited hycean scenarios. We show in Figure 12 that, even if one assumes a hycean scenario with a large unknown CH4 flux (e.g. from methane-producing life), over time the atmospheric CO2 and CO abundances would build to levels comparable to or above our upper limits. Consequently, our non-detections of CO2 and CO are not consistent with the observed CH4 abundance under an inhabited hycean scenario. Our results therefore render the mini-Neptune scenario the most likely interpretation for K2-18 b given current observational constraints.
@@
Our revised atmospheric composition also motivates a careful reevaluation of the viability of methanogenesis on K2-18 b. Our non-detections of CO2 and DMS prompt a re-assessment of whether a sufficient free energy source exists for methane-producing life under the metabolic pathway studied by Glein (2024), rendering it less plausible that K2-18 b sustains a biosphere. Further, our models in Figure 12 demonstrate that a hycean scenario with the CH4 abundance we retrieve would necessarily have a high 1% CO2 abundance that is in slight tension with our non-detection of CO2 (though it is consistent with our 2- upper limit). Updated metabolic calculations that consider the minimum viable CO2 abundance to sustain methanogenesis will be required to assess whether a biosphere is ruled out by our revised atmospheric composition.